sum of three consecutive odd no. is 63.find them
Answer by Aswathy(23)

Let 3 consecutive odd numbers be 2x+1,2x+3 and 2x+5.
Then according to the given conditions,
(2x+1)+(2x+3)+(2x+5)=63
6x+9=63
6x=54
x=54/6
x=9
2x+1=19
2x+3=21
2x+5=23
Thus the three consecutive odd numbers are 19,21 & 23.